Telephony Seminar

6.913

Final Report

André DeHon

This semester 6.913 followed the progression of basic telephone technology from the invention of the telephone by professor Bell through a number of the current electronic switching systems in use today. The technology was surveyed both from the point of view of the individual receivers and that of the switching systems. Throughout the course, in addition to reviewing the electronic innovation involved in each switching systems, the manner in which each system met the needs of arbitration, control, and accounting was emphasized.

The telephone had its origins in telegraph research. The concept of actually transmitting voice across wires in this manner began with professor Alexander Bell. While he was attempting to send multiple signals down a telegraph line, he noted that he could get a tone at one end of a line to actuate a resonate structure at the other. With this he began using various structures, such as diaphragms and aciduated water, to convert acoustic signals into voltages which would then be transmitted across a wire and decoded at a receiver.

With the advent of such a means of transmitting sound over distances, people began to make use of the new technology. Initially, a number of “telephones” were connected to each other via a common telegraph-like wire thus providing a primitive party line for those who shared the connection. Such early systems were very ad hoc, and thus not regulated. Neither did they have any real means of switching, control, or routing. Eventually, as a first attempt to deal with the issue of signalling over a phone system, a crank was added to wiggle the transducer on a phone in a characteristic manner in order to both inform subscribers that a call is being requested and to distinguish the intended recipient.

Soon, the crude coupling provided by directly connecting transducers was improved upon. A second line was introduced to provide a common point of reference for transmitted voltage signals. The batteries within each phone were decoupled using first capacitors, then transformers. Early on, a separate bell was added to indicate that a call was pending for a particular locale. The functions of the receiver and transmitter were split into two distinct pieces of hardware. Much improvement in the quality of signal transmitted occurred with the introduction of the carbon microphone as a common base for converting acoustic signals such as speech into electrical currents. Also the receiver and the transmitter were eventually decoupled via transformers in order to minimize side tones and feedback.

As the number of phones in user grew, more and more connectivity was desired. Originally, the desired party was indicated by the number of rings in the ringing sequence. Connecting everything point to point or even as large party-lines was clearly unreasonable; additionally, as the number of subscribers on a party-line increased, it both became difficult to distinguish the destination specified and for an individual to get a reasonable opportunity to use the phone when he wanted. As a result, a central switching scheme was devised to allow connections between local party lines and more distant locations.

To meet this need, switching systems came into being around 1880. The first had electrical connections for each line the system serviced. Then it had a fixed number of selectors which could be used to route a call from an originator to a destination. The switching, routing control, and signalling were virtually all supplied by a human attendant. The attendant would be signalled that a subscriber wished to route a call. From there, the attendant would attempt to satisfy the subscribers request by signalling the destination addressee and routing the call when possible. This switching scheme proved to be quite limited in the number of calls that could be connected at a given time.

The next generation of switching was the plug and jack switch board. This still relied on a human attendant for most of the routing and arbitration, but provided a more sophisticated set of hardware tools for the attendant to work with. Each line terminated in jack which represented the important parts of a given line (the tip and sleeve). Each line also had an additional plug that could be used to connect a jack from another line to. This allowed the arbitrary connection of one line to another without limiting the number of calls that could be routed simultaneously as severely as the earlier generation. This system provided the attendant with a direct way to test if the line was in use, provide ringing request for attention to a subscriber, and obtain notification that a call was terminated (via ring off). Along with this generation, switch boards such as this also incorporated a number of lines to other switch boards so that connections could be made to subscribers outside of the immediate population serviced by the given switchboard.

The next generation of switching systems introduced more automated signalling and routing. This was initiated with the step by step switching system. Here the idea was to convey routing information along with the call and provide a progressive automated switching. Also by this time the battery to provide power for the phone conversations was moved from the individual hand-sets to the central switching office. The step by step system was then able to note a request for service from a local phone set by checking the current drawn by the local set. When a request for service is detected, the switching system attempts to find a free line to use for the call. Once a line is found, the subscriber may “dial” the desired destination. The dial on the hand set produces a sequence of pulses. The step by step system uses each sequence to specify a routing connection through the office. This is done by stroger switches which use each pulse to advance a mechanical arm to a different pad to route the connection. In this manner, each “digit” in the number is associated with a different piece of equipment. The line finders concentrate the incoming calls and respond to the service request. The incremental switching through stroger style switches provides distribution, routing, and expansion to connections with other phones in or out of the office. On top of this, special offices called tandem offices were available to provide indirect routing from point to point. If a given switchboard couldn’t make a direct connection to a destination, it would then route the call through a tandem office.

While the step by step automated much of the arbitration of call routing, it had a number of draw backs. The number of connections that could be made was limited by the number of line finders and number switches for given digits. Also the system was very static. The system was difficult to reconfigure with changing usage patterns. Also, it required physical reconfigurations to change a telephone number.

The next generation, the panel office sought to remedy some of these deficiencies. In the panel office, the serviced lines were arranged in a switch network. The panel switch had a traveling vertical contact which looked for active lines. For the most part, the panel office still used the sequences of pulses for specifying the desired user routing. When the digits were transmitted by the local hand set, a component known as the sender would store the digits in a special relay, decode the stored value, and translate this into the number of pulses necessary to set up the connection in the switch matrix.^{1} In this system all of the information necessary to connect the call was resident in the machine rather than being distributed over several switches as in the step by step scheme. Additionally, this scheme provided more direct means of connecting to trunks which allowed connections to other switching offices for requests for service to numbers not directly serviced by the local switch. This matrix scheme provided a much more uniform means of connection without the necessity of dedicating multiple stroger switches to each call. Also it allowed for the more dynamic reconfiguration missing in the step by step system.

1This operation relied on feedback from the panel switch in order to position the connection accurately.

The panel offices were eventually superseded by the crossbar offices. The crossbar theory was similar to that of the panel office, but it provided a different means of implementation. A line finder was used to find connect a line requesting service to an open line in the cross bar switching. A Marker would then then link the incoming lines via several crossbar frames first to lines to the trunks, then to the respective trunks out of the office or to the appropriate intra-office connections. Additionally, a crossbar frame services incoming calls from outside of the office to connect these corresponding lines to the requested local subscriber lines. The crossbar office proved to be much more maintainable than its predecessors. Also, the crossbar office was designed to provide more automatic accounting of system usage by subscribers. The crossbar office expands on the theme of reconfigurability for growth and change like the panel office.

The next generation of switching was the electronic office. Here solidstate electronic devices began to replace the electromechanical equipment used in earlier systems. Utilizing the more advanced technology, this allowed more versatility, faster operation, easier maintenance, and more services for the subscriber. The electronic office provides semipermanent programmable memory that defines the current configuration of the system. This includes the translation of a number to a physical wire, what services a subscriber has, and other fixed information about a given subscriber. The office also has a dynamic memory to keep track of current call status (e.g. which phones are busy/free, dialing information, billing information). Permanent memory (e.g. magnetic tape) is used to actually store billing information for completed calls. The individual components are modular mechanical devices and easily replaced when problems arise; the system diagnostics should be capable of identifying faulty equipment which can be replaced as necessary to keep equipment in operation. While the technology is much more advanced with the electronic office than the earlier offices, the switching scheme itself is somewhat similar. Lines into the office are polled for service requests.

These are then linked through junctors to the appropriate electronic routing to establish either a local talk path, or a path through the trunks to remote switching to reach a location outside of the local office. In the electronic office, the operation of the entire switching system is moderated by high speed computer which uses the semipermanent memory as a pseudo-program and the various lines and switching as its peripheral.

Also of importance in the development of modern telephone technology is the transmission of the call information over wires. Conversations over telephone wires suffer from attenuation, dispersion,^{2} and echoing. The distortion and attenuation effects were ameliorated by carefully loading the transmission line. The loss in signal strength over distance in long distance lines was first dealt with using repeaters. As opposed to telegraphs, phones required a second line so that random noise in the line could be eliminated by observing only the differential signal on the pair of lines. Eventually, in order to multiplex multiple conversations over a single line pair, modulation was used as over long distance lines. In this manner, a number of conversations shared a line at different frequencies. The multiplexed scheme also reduced dispersion as the band width was smaller relative to the transmission frequency.

Recently, digital sampling has come to successful usage in telephone line transmission. Using sampling, the local office can sample the call at 8kHz. Then it can use its local memory to store the sampled speech data and time multiplex it over an outgoing long distance line with an additional bit of routing information. In this manner, multiple conversations can be conveyed over a signal high band width data line digitally by sending pieces of each conversation interleaved and reconstructing the digital information into an audio signal at the other end. This also has the advantage that the digital information does not suffer attenuation, dispersion, or echoing as the analog signal had. At each step, the digital information can be checked using some check-sum/parity scheme to see that it is correct and an update requested if it is not.

2delay distortion